# PATCH - Psychometrics-AssisTed benCHmarking of Large Language Models: A Case Study of Mathematics Proficiency

### v1

@@ 3.3 LLM Benchmark Development
Developing LLM benchmarks follows a similar yet different process. Take GSM8K (Cobbe et al., 2021) as an example. According to the GSM8k paper, the authors started by specifying the need for a large, high quality math test at grade school level that is of moderate difficulty for LLMs (Step 1). The implied construct (PGSM) is not linked to any specific curriculum. Then, the overall planning is made (Step 2): The number of items should be in the thousands; The items will be curated by crowd workers; Agreement and error analysis will be used to investigate the quality of the dataset; GPT-3 will be used to benchmark the dataset and verify the difficulty of the dataset; etc. In Step 3, where dataset development takes place, often one of the two strategies is used: either collect items from existing datasets and other sources and compile them into a new dataset, or (like GSM8K) create own items from scratch possibly with annotations. The latter is usually an iterative procedure consisting of four parts: creating instructions (and possibly a user interface) for item generation and/or annotation; conduct a (pilot) study to collect the items and/or annotations; check annotator agreement; and assess errors associated with the items or annotations. This step is iterated until a sufficient number of items is reached that meets desired quality standards (e.g., high annotator agreement, low error rate). In total, GSM8K collects 8,500 items with solutions, with identified annotator disagreements resolved and a less than 2% error rate. In Step 4, the generated items form the final dataset, which is typically split into training, evaluation and testing partitions. In Step 5, the final selection of LLMs is made and evaluated on the dataset. Finally, in Step 6, the benchmark gets released, which typically consists of the dataset as well as its documentation (often a research paper) and benchmarking results.
@@
While sharing similarity with test development in psychometrics, benchmark development for LLMs falls short on four aspects. First, the construct of interest is often under-specified, leading to a mismatch between the intended construct and what the dataset actually measures. Take GSM8K as an example: while the dataset is intended to measure proficiency in grade school mathematics, the target grade level(s) are unclear and it only focuses on one content domain (algebra), missing other relevant ones like geometry and data. This is likely the result of not using established mathematics curriculum and domain experts to develop test items. Second, despite LLM researchers’ interest in comparing LLM performance with human test takers (e.g., the GSM8K paper claims that “a bright middle school student should be able to solve every problem”), such comparisons usually cannot be made because the test has not been designed with humans in mind or validated on any (representative samples of) human test taker populations. Third, in addition to agreement and error analysis, LLM benchmarks can benefit from psychometric analysis of test items, (i.e., checking item discriminativeness and difficulty, as well as the factor structure of the items). While this is currently not the norm, there have been some promising attempts (see § 2). Lastly, the released benchmark often does not contain sufficient details about all the steps involved in creating the benchmark. For instance, the GSM8K paper does not include the instructions for item creation and annotation, the results from the pilot study, the agreement numbers, or annotator characteristics, all of which are important for external researchers to independently validate the quality of the benchmark.
@@ 4 PATCH: Psychometrics-Assisted benCHmarking of LLMs
Figure 1 illustrates PATCH, our conceptualization of a psychometrics-assisted framework for benchmarking LLMs [^5] Partly inspired by the Hexagon Framework of scientific measurements in Mari et al. (2023).. In PATCH, the first step is to define the construct of interest (e.g., proficiency in 8th grade mathematics). The second step is to look for an existing validated psychometric test that measures this property; alternatively, a test can be developed from scratch, following the procedures described in § 3.2, which likely requires collaboration with experienced psychometricians. The term ”validated” means that the test has been tested on a representative sample of the target population of human test takers and fulfills several psychometric quality requirements (e.g., discriminative items that are well distributed across different difficulty levels; showing high reliability (e.g., high internal consistency) and validity (e.g., the test’s factor structure matches the construct definition)).
@@
Next (Step 3$\rightarrow$4), we use the items in the validated psychometric test to construct prompts for the LLMs under evaluation, and then sample responses. A response typically consists of a task description, an explanation and an answer (key). Therefore, in Step 4$\rightarrow$5, we extract the answer (key) for each item’s response, then grade it to obtain item scores (Step 5$\rightarrow$6).
For Step 2$\rightarrow$7, the responses of human test takers can be used to estimate IRT item parameters and subsequently the latent proficiency scores for each test taker (human or LLM) with uncertainty estimates. Multiple IRT models are often used because of the use of different types of test items. These latent scores are typically standardized $z$-scores (i.e., mean of 00 and SD of $1$), which can optionally go through further normalization (e.g., re-scaling to a mean of $500$ and a SD of $100$) (Step 6$\rightarrow$7). These final proficiency scores can be used for comparison with other models and populations.
It can be said that the heart of PATCH is a validated psychometric test, which not only provides the basis for accurate measurement of a model’s capability of interest but also facilitates comparison between LLMs and human test takers. Unfortunately, developing such a test can be a long and expensive process; utilizing existing tests can be a shortcut, which, however, need to satisfy three requirements: clear human population reference; test items available (released); human responses and/or item parameter estimates available. The second and third requirement are in practice difficult to meet, as many test institutes do not make their test items public due to commercial interests (e.g., SAT) or the need to measure trends over time (e.g., PISA). Collaboration with test institutes would be ideal.
To the best of our knowledge, when it comes to academic proficiency tests, only TIMSS and PIRLS tests from certain years can be readily used for PATCH-based LLM benchmarking. TIMSS measures proficiency in grade school mathematics and science (4th grade, 8th grade, and final year of secondary school), while PIRLS assesses reading comprehension in 9-10 year olds. Both TIMSS and PIRLS are administered in a large number of countries and regions with representative student samples, enabling country/region-level comparisons. In the following section, we demonstrate PATCH by measuring GPT-4 and Gemini’s proficiency in 8th grade mathematics, using the latest available data from TIMSS: TIMSS 2011.
@@ 5.1 Data: TIMSS 2011 8th Grade Mathematics
The TIMSS 2011 study has released 88 mathematics test items, 48 of which are multiple choice, 30 open-ended items scored as either incorrect or correct, and 10 open-ended items scored as either incorrect, partially correct, or correct. These items assess four content domains representative of 8th grade mathematics curriculum (agreed upon by experts from participating countries/regions): number, algebra, geometry, data and chance. Within each domain, items are designed to cover various subtopics (e.g., decimals, functions, patterns) and three cognitive domains: knowing, applying and reasoning. These test items are only available in a PDF file that can be downloaded from the NCES website, which includes also scoring instructions. [^7] https://nces.ed.gov/timss/pdf/TIMSS2011_G8_Math.pdf To extracting them into a format compatible with LLMs, we used OCR tools to extract as much textual information as possible, converted mathematical objects (e.g., numbers, symbols, equations, tables) into LaTeX format (following earlier benchmarks like MATH (Hendrycks et al., 2021)) and figures into JPEG format. See Appendix A.1 for examples. We will release this LLM-compatible version of test items, as well as a mathematics dataset from TIMSS 2008 and two science datasets from TIMSS 2011 and 2008.
@@
The second part of the dataset concerns the specific IRT model used for each test item and the estimated item parameters (e.g., discriminativeness, difficulty), which can be used to reconstruct the IRT formulas for estimating proficiency scores.
@@
Lastly, the students’ responses to each test item and their proficiency estimates are available, allowing us to construct proficiency score distributions for each country and region.
@@ 5.2 LLMs: GPT-4 with Vision and Gemini-Pro-Vision
A major concern in using these closed-source LLMs is data contamination, which is difficulty to check due to inaccessible training data. However, as our focus is on demonstrating the PATCH framework, data contamination is less worrying. Furthermore, data contamination is still unlikely for four reasons. First, these test items are copyrighted, forbidding commercial use. Second, the test items are hard to extract from the PDF file. Third, to the best of our knowledge, these test items do not exist in current LLM mathematics benchmarks. Fourth, we request GPT-4V and Gemini-Pro-Vision to explain or provide solutions to the test items’ IDs (available in the PDF file). Both failed to recognize these specific test IDs.
@@ 5.3 Prompts and Temperature
We design two separate prompts for each test item: the system message and the user message. We design the system message according to the prompt engineering guide by OpenAI [^8] https://platform.openai.com/docs/guides/prompt-engineering, utilizing chain-of-thought and step-by-step instructions on how to respond to the user message (i.e., with a classification of question type, an explanation and an answer (key)). The system message is the same for all test items (see Appendix A.2). Furthermore, to account for LLMs’ sensitivity to slight variations in prompts (e.g., Sclar et al., 2023; Loya et al., 2023), we generate 10 additional variants of the system prompt with slight perturbations (e.g., lowercase a heading, vary the order of unordered bullet points).
@@ 5.4 Response Scoring and Proficiency Estimation
We manually examine the sampled responses from GPT-4V and Gemini-V and score them following the official scoring guide of TIMSS 2011. Then, for multiple choice items, we apply the 3PL model (Equation 1); for open-ended items, we apply the GPC model (Equation 2) if partially correct response is admissible, otherwise the 2PL model. We use maximum likelihood to obtain unbiased estimates of model proficiency scores ($\theta$) with the mirt package in R (Chalmers, 2012). This results in 11 $\theta$ estimates per model due to the use of 11 system message variants. We then use inverse variance weighting (Marín-Martínez & Sánchez-Meca, 2010) to obtain a weighted $\theta$ and its 95% confidence interval (CI) for each model.
@@ 5.5 Results
Figure 2: Distribution of proficiency estimates for GPT-4V, Gemini-Vision-Pro and selected participants of the TIMSS 2011 8th grade mathematics test. Left figure (A) shows the proficiency estimates based on the percentages of correct responses. Right figure (B) shows the IRT-based proficiency estimates. The middle vertical line in each box plot represents the weighted mean proficiency score, with the error bars indicating its 95% confidence interval. The borders of each box indicate the range of the middle 50% of all values, with the two whiskers indicating the 5th and 95th percentiles.
Figure 2 shows the proficiency score distribution and ranking of 15 selected participating countries & regions, GPT-4V and Gemini-Pro-Vision. [^10] Only 15 countries are shown here to save space. The complete figures can be found in Appendix C. The proficiency scores on the left panel (x-axis) are percentages of correct responses, which is the default approach in current LLM benchmarking; the proficiency estimates on the right panel are based on IRT. We make three observations. First, regardless of the method of proficiency estimation, GPT-4V has the overall best performance relative to Gemini-Pro-Vision and the average proficiency of 8th grade students of each participating country/region. Second, the method of proficiency estimation affects the ranking results. For instance, while Chinese Taipei is ranked third on the left, it is ranked 4th on the right; Gemini-Pro-Vision is ranked the 8th on the left, but ranked the 7th on the right. Third, the method of proficiency estimation affects the estimated 95% CIs, which are usually wider when IRT is used (as it accounts for both item and test taker variances). Notably, while on the left panel the CI of GPT-4V does not overlap with the second best, South Korea, indicating a statistically significant difference, they overlap on the right panel, suggesting otherwise. This finding shows that the adoption of PATCH is likely going to make a difference to LLM benchmark results.
@@ 6 Conclusions
In this paper, we propose PATCH, a psychometrics-inspired framework to address current limitations of LLM benchmarks, especially for the purpose of model and human comparison. We demonstrate PATCH with a 8th grade mathematics proficiency test, where PATCH yields evaluation outcomes that diverge from those based on existing benchmarking practices. This underscores the potential of PATCH to reshape the LLM benchmarking landscape. Nevertheless, our paper has several limitations. First, PATCH requires validated tests, which can be resource-intensive if tests need to be developed from scratch. However, this also opens up opportunities for collaboration between LLM researchers, psychometricians and test institutes. Second, the validity, reliability, and fairness of using tests validated solely on humans for LLM benchmarking are debatable due to possibly differing notions of proficiency and cognitive processes between LLMs and humans. Nonetheless, such tests are still better than non-validated benchmarks, particularly for comparison of model and human performance. Advancing LLM benchmarking further requires tests validated on LLMs (and humans for model-human comparisons), necessitating theoretical work on LLM-specific constructs and the development of LLM-specific IRT models and testing procedures.
@@ A.2 Example System Messages
You are given a math question written in LaTeX format.

### v2

@@ 3.3 LLM Benchmark Development
Developing LLM benchmarks follows a similar yet different process. Take GSM8K (Cobbe et al., 2021) as an example. According to the GSM8k paper, the authors started by specifying the need for a large, high quality mathematics test at grade school level that is of moderate difficulty for LLMs (Step 1). The implied construct (i.e., PGSM) is, however, not explicitly linked to any specific curriculum.
Then, the overall planning is made (Step 2): The number of items should be in the thousands; the items will be curated by crowd workers; agreement and error analysis will be used to investigate the quality of the dataset; GPT-3 will be used to benchmark the dataset and verify the difficulty of the dataset.
In Step 3, where dataset development [^8] Note that we use the term “dataset development” here, contrasting “item development” in psychometrics, because of LLM benchmarks’ typical emphasis on large and multiple datasets rather than concrete test items. takes place, often one of the two strategies is used: either collect items from existing datasets and other sources and compile them into a new dataset, or, like in GSM8K, create own items from scratch (with annotations). The latter is usually an iterative procedure consisting of four parts: creating instructions (and possibly a user interface) for item generation and/or annotation; conduct a (pilot) study to collect the items and/or annotations; check annotator agreement; and assess errors associated with the items or annotations. This step is iterated until a sufficient number of items and datasets is reached while meeting desired quality standards (e.g., high annotator agreement, low error rate). In total, GSM8K includes 8,500 items with solutions, with identified annotator disagreements resolved and a less than 2% error rate.
In Step 4, the generated items make up the final dataset, typically split into training, evaluation and testing partitions. In Step 5, the final selection of LLMs is made and evaluated on the dataset. Finally, in Step 6, the benchmark gets released, which typically consists of the dataset as well as its documentation (often a research paper) and benchmarking results.
@@
While sharing similarity with test development in psychometrics, benchmark development for LLMs falls short on four aspects. First, the construct of interest is often under-specified, leading to a mismatch between the intended construct and what the dataset actually measures. Take GSM8K as an example: While the dataset is intended to measure proficiency in grade school mathematics, the target grade level(s) are unclear and it only focuses on one content domain (algebra), missing other relevant ones like geometry and data. This is likely the result of not using established mathematics curricula and domain experts to develop test items.
Second, despite researchers’ interest in comparing LLM performance with human test takers (e.g., the GSM8K paper claims that “a bright middle school student should be able to solve every problem”), such comparisons usually cannot be made because the test has not been designed with humans in mind or validated on any representative samples of the test’s target user populations.
Third, besides agreement and error analysis, LLM benchmarks can benefit from psychometric analysis of test items, (i.e., checking item discriminativeness and difficulty, as well as the factor structure of the items). While this is not yet the norm, there have been promising attempts (see Section 2).
Lastly, the released benchmark often does not contain sufficient details about all the steps involved in creating the benchmark. For instance, the GSM8K paper does not present the instructions for item creation and annotation, the results from the pilot study, the agreement statistics, or annotator characteristics, all of which are important for external researchers to independently validate the quality of the benchmark.
@@ 4 PATCH: Psychometrics-AssisTed benCHmarking of LLMs
Figure 1 illustrates PATCH, our conceptualisation of a psychometrics-assisted framework for benchmarking LLMs. [^9] PATCH is partly inspired by the Hexagon Framework of scientific measurements proposed by Mari et al. (2023). According to PATCH, the first step is to define the construct of interest (e.g., proficiency in 8th grade mathematics).
The second step is to look for an existing validated psychometric test that measures this property; alternatively, a test can be developed from scratch, following the procedures described in Section 3.2, which likely requires collaboration with experienced psychometricians. The term “validated” means that the test has been tested on a representative sample of the target population of human test takers and fulfils several psychometric quality requirements (e.g., discriminative items that are well distributed across different difficulty levels; showing high reliability (e.g., high internal consistency) and validity (e.g., the test’s factor structure matches the construct definition)).
@@
Next (Step 3$\rightarrow$4), we use the items in the validated psychometric test to construct prompts for the LLMs under evaluation and then sample responses. A response typically consists of a task description, an explanation and an answer (key). Therefore, in Step 4$\rightarrow$5, we extract the answer (key) for each item’s response, then grade it to obtain item scores (Step 5$\rightarrow$6).
For Step 2$\rightarrow$7, the responses of human test takers (and of LLMs, if a sufficient number of LLMs are involved) can be used to estimate IRT item parameters and subsequently the latent proficiency scores for each test taker (human or LLM) with uncertainty estimates. Multiple IRT models are often used because of the use of different types of test items. These latent scores are typically standardised $z$-scores (i.e., mean of 00 and standard deviation of $1$), which can optionally go through further normalisation (e.g., re-scaling to a mean of $500$ and a standard deviation of $100$) (Step 6$\rightarrow$7). These final proficiency scores can be used for comparison with other models and populations.
It can be said that the heart of PATCH is a validated psychometric test, which not only provides the basis for accurate measurement of a model’s capability of interest but also facilitates comparison between LLMs and human test takers. Unfortunately, developing such a test can be a long and expensive process; utilising existing tests can be a shortcut, which, however, should satisfy three requirements: clear human population reference; test items available (released); human responses and/or item parameter estimates available. The second and third requirement are in practice difficult to meet, as many test institutes do not make their test items public due to commercial interests (e.g., SAT) or the need to measure trends over time (e.g., PISA). Collaboration with test institutes would alleviate this problem.
To the best of our knowledge, when it comes to academic proficiency tests, only TIMSS and PIRLS tests from certain years can be readily used for PATCH-based LLM benchmarking. TIMSS measures proficiency in grade school mathematics and science (4th grade, 8th grade, and final year of secondary school), while PIRLS assesses reading comprehension in 9/10-year-olds. Both TIMSS and PIRLS are administered in a large number of countries and regions with representative student samples, enabling country/region-level comparisons. In the following section, we demonstrate PATCH by measuring GPT-4 and Gemini’s proficiency in 8th grade mathematics, using the latest available data from TIMSS 2011.
@@ 5.1 Data: TIMSS 2011 8th Grade Mathematics
The TIMSS 2011 study has released 88 mathematics test items, 48 of which are multiple choice, 30 open-ended items scored as either incorrect or correct, and 10 open-ended items scored as either incorrect, partially correct, or correct. These items assess four content domains representative of 8th grade mathematics curriculum (agreed upon by experts from participating countries/regions): number, algebra, geometry, data and chance. Within each domain, items are designed to cover various subtopics (e.g., decimals, functions, patterns) and three cognitive domains: knowing, applying and reasoning. These test items are only available in a PDF file that can be downloaded from the NCES website, which includes also scoring instructions. [^11] https://nces.ed.gov/timss/pdf/TIMSS2011_G8_Math.pdf To extracting them into a format compatible with LLMs, we used OCR tools to extract as much textual information as possible, converted mathematical objects (e.g., numbers, symbols, equations, tables) into LaTeX format (following earlier benchmarks like MATH (Hendrycks et al., 2021)) and figures into JPEG format. See Appendix A.1 for examples. We have released this LLM-compatible version of test items, as well as an eighth grade science test dataset from TIMSS 2011, an advanced secondary school mathematics test dataset from TIMSS 2008, and an advanced secondary school physics test dataset from TIMSS 2008121212Available via https://zenodo.org/records/12531906.
@@
The second part of the dataset concerns the specific IRT model used for each test item and the estimated item parameters (e.g., discriminativeness, difficulty), which can be used to reconstruct the IRT models for estimating proficiency scores.
@@
Lastly, responses of the sampled students (about 4,000 on average per country/region) to each test item and their proficiency estimates have also been made available, allowing us to construct proficiency score distributions for each country and region.
@@ 5.2 LLMs: GPT-4 with Vision and Gemini-Pro-Vision
A major concern in using these closed-source LLMs is data contamination, which is difficulty to check due to inaccessible (information about) training data. However, as our focus is on demonstrating the PATCH framework, data contamination is less worrying. Furthermore, data contamination is still unlikely for four reasons. First, these test items are copyrighted, forbidding commercial use. Second, the test items are hard to extract from the PDF file. Third, to the best of our knowledge, these test items do not exist in current LLM mathematics benchmarks. Fourth, we ask GPT-4V and Gemini-Pro-Vision to explain or provide solutions to the test items’ IDs (available in the PDF file). Both failed to recognise these specific test IDs.
@@ 5.3 Prompts and Temperature
We design two separate prompts for each test item: the system message and the user message. We design the system message according to the prompt engineering guide by OpenAI, utilising chain-of-thought and step-by-step instructions on how to respond to the user message (i.e., with a classification of question type, an explanation and an answer (key)). [^13] https://platform.openai.com/docs/guides/prompt-engineering The system message is the same for all test items (see Appendix A.2). Furthermore, to account for LLMs’ sensitivity to slight variations in prompts (e.g., Sclar et al., 2024; Loya et al., 2023), we generate 10 additional variants of the system prompt with slight perturbations (e.g., lowercase a heading, vary the order of unordered bullet points).
@@ 5.4 Response Scoring and Proficiency Estimation
We manually examine the sampled responses from GPT-4V and Gemini-V and score them following the official scoring rubrics of TIMSS 2011. Then, for multiple choice items, we apply the 3PL model (Equation 1); for open-ended items, we apply the GPC model (Equation 2) if partially correct response is admissible, otherwise the 2PL model. We use maximum likelihood to obtain unbiased estimates of model proficiency scores ($\theta$) with the mirt package in R (Chalmers, 2012). This results in 11 $\theta$ estimates per model due to the use of 11 system message variants. We then use inverse variance weighting (Marín-Martínez & Sánchez-Meca, 2010) to combine these estimates. Inverse variance weighting gives more weight to estimates that are more precise (i.e., having lower variance) and less weight to those that are less precise (i.e., having higher variance). This way, we obtain a more accurate overall $\theta$ estimate and its 95% confidence interval (CI) for each model.
@@ 5.5 Results
Figure 2: Distribution of proficiency estimates for GPT-4V, Gemini-Vision-Pro and selected participating countries/regions of the TIMSS 2011 8th grade mathematics test. Left figure (A) shows the proficiency estimates based on the percentages of correct responses. Right figure (B) shows the IRT-based proficiency estimates. The middle vertical line in each box plot represents the weighted mean proficiency score, with the error bars indicating its 95% confidence interval. The borders of each box indicate the range of the middle 50% of all values, with the two whiskers indicating the 5th and 95th percentiles.
Figure 2 shows the proficiency score distribution and ranking of 15 selected participating countries and regions, GPT-4V and Gemini-Pro-Vision. Only 15 countries are shown here to save space. The complete figures can be found in Appendix C. The proficiency scores (x-axis) on the left panel are percentages of correct responses, which is the default approach in current LLM benchmarking; the proficiency estimates on the right panel are based on IRT. We make three observations. First, regardless of the method of proficiency estimation, GPT-4V has the overall best performance relative to Gemini-Pro-Vision and the average proficiency of 8th grade students of each participating country/region. Second, the method of proficiency estimation affects the ranking results. For instance, while Chinese Taipei is ranked 3rd on the left, it is ranked 4th on the right; Gemini-Pro-Vision is ranked 8th on the left, but ranked 7th on the right. Similarly, while Hungary is ranked 11th on the left, it drops to the 16th place on the right. Third, the method of proficiency estimation affects the estimated 95% CIs, which are usually wider when IRT is used (as it accounts for both item and test taker variances). Notably, while on the left panel the CI of GPT-4V does not overlap with the second best, South Korea, indicating a statistically significant difference, they overlap on the right panel, suggesting otherwise. This finding shows that the adoption of PATCH is likely going to make a difference to LLM benchmark results.
@@ 6 Conclusions
In this paper, we propose PATCH, a psychometrics-inspired framework to address current limitations of LLM benchmarks, especially for the purpose of model and human comparison. We demonstrate PATCH with an 8th grade mathematics proficiency test, where PATCH yields evaluation outcomes that diverge from those based on existing benchmarking practices. This underscores the potential of PATCH to reshape the LLM benchmarking landscape. Nevertheless, our paper has several limitations. First, PATCH requires validated tests, which can be resource-intensive if tests need to be developed from scratch. However, this also opens up opportunities for collaboration between LLM researchers, psychometricians and test institutes. Second, the validity, reliability, and fairness of using tests validated solely on humans for LLM benchmarking are debatable due to possibly differing notions of proficiency and cognitive processes between LLMs and humans. Nonetheless, such tests are still better than non-validated benchmarks, particularly for comparison of model and human performance. Advancing LLM benchmarking further requires tests validated on LLMs (and humans for model-human comparisons), necessitating theoretical work on LLM-specific constructs and the development of LLM-specific IRT models and testing procedures. Third, our experiment only includes two proprietary LLMs and one proficiency test. We consider this sufficient for demonstrating PATCH, but not enough if the goal is to benchmark all relevant LLMs across different tests.
Acknowledgements
We thank Anna Wegmann, Yupei Du, Melody Sepahpour-Fard, Elise Herrewijnen, Gianluca Sperduti for their helpful suggestions and comments. This work was supported by the Dutch Research Council (NWO) (grant number VI.Vidi.195.152 to D. L. Oberski; grant number VI.Veni.192.130 to D. Nguyen).
@@ A.2 Example System Messages
You are given a mathematics question written in LaTeX format.